# Multi-Scale Data Assimilation in Turbulent Models

### v1

@@ Abstract
We explore the potential of Data-Assimilation (DA) within the multi-scale framework of a shell model of turbulence, with a focus on the Ensemble Kalman Filter (EnKF). The central objective is to understand how measuring mesoscales (i.e., inertial-range scales) enhances the prediction of both large-scale and small-scale intermittent variables, by systematically varying observation frequency and the set of measured scales. We demonstrate that measurements conducted at frequencies that exceed those of the observed scales enable full synchronization of larger scales, provided that at least two adjacent mesoscale are measured. In addition, we benchmark the EnKF against two standard DA methods—Nudging and Ensemble 4D-Var—showing its overall superior performance. Moreover, our results underscore the need for a tailored, scale-aware inflation technique to stabilize the assimilation process, preventing filter divergence and ensuring robust convergence.
@@ I Introduction
In order to overcome these limitations, the Data Assimilation (DA) framework was developed. Originally motivated by meteorological forecasting, DA provides systematic methods to optimally combine model forecasts with sparse and noisy observations, improving the estimation of system states [6, 7, 8]. In chaotic systems, DA offers a practical strategy to counteract the exponential divergence of trajectories by acting as an external forcing—analogous to a drag—that competes with internal instabilities [9]. This competition naturally defines a synchronization threshold: When observational corrections outweigh the intrinsic divergence rate, synchronization can be achieved, and thus effective predictability.
@@
This work mainly focuses on investigating the performance of a cornerstone DA technique, the Ensemble Kalman Filter (EnKF) [10, 11], when applied to a shell model for turbulence [12, 13, 14, 15]. Our objective is to reconstruct the unresolved scales from sparse and noisy observations of the resolved ones. Shell models emulate the turbulent energy cascade in a discretized Fourier space with logarithmically spaced wavenumbers, yielding several decades of scales and an equally wide spectrum of characteristic turnover times (Fig. 1(a)). Assimilating information measured at one scale to correct predictions at another is therefore an intrinsically multiscale challenge: updates must bridge slow, energy-containing modes and fast, highly intermittent ones.
@@
The EnKF advances an ensemble of model realizations and, at every assimilation instant, applies statistically optimal linear updates, provided the forecast and observation errors are Gaussian. Because the ensemble itself furnishes flow-dependent (prior) error covariances, the EnKF dispenses with the tangent-linear and adjoint operators demanded by the classic Kalman filter. However, in high-dimensional settings, computational constraints limit the ensemble size, introducing sampling noise that can erode covariance rank and trigger filter divergence [7, 16]. Practitioners therefore resort to inflation and covariance localization to maintain the need of ensemble spread and numerical stability [17, 18, 19, 20]. Despite these practical remedies, the EnKF and its extensions remain among the most widely used and readily implementable data assimilation methods for large-scale geophysical problems [8, 16]. These limitations become particularly important in the multiscale, strongly intermittent dynamics of turbulence. The shell model considered here spans a few decades of scales and turnover times, and it faithfully reproduces the heavy-tailed, non-Gaussian statistics characteristic of small-scale turbulent fluctuations (see Fig. 1(c)). Such departures from Gaussianity degrade the nominal quasi-optimality of deterministic EnKF variants. For this reason, we adopt the stochastic EnKF, which injects controlled Gaussian perturbations into the observations—or equivalently, into each ensemble member—at every analysis step. This added noise mitigates sampling error, enhances ensemble dispersion, and partly compensates for the non-Gaussian heavy tails generated by the fastest, most intermittent scales, thereby improving filter robustness in the highly nonlinear, multiscale regime addressed here. The present contribution therefore constitutes a first step toward systematically tailoring ensemble-based DA techniques to highly turbulent flows.
@@
We shall benchmark the EnKF against two Data Assimilation methods at opposite ends of the DA complexity hierarchy. First, Nudging [21], which can be viewed as a stripped-down single-member analogue of EnKF. By adding continuous relaxation terms toward observations, it requires almost no statistical assumptions and is computationally inexpensive, but its strictly local corrections struggle to propagate information to unmeasured scales. The propagation of information from observations throughout the full domain is done solely via the model dynamics, and, as opposed to the EnKF it does not provide a quantification of the uncertainty in the state estimate.
@@
Second, we consider the so-called four‐dimensional variational method (4D-Var) [22]. 4D-Var seeks the model trajectory that minimizes a time-integrated cost function over a finite assimilation window, thereby distributing corrections smoothly in time. The 4D-Var is formulated as a nonlinear optimization problem, as opposed to the linear-like update in the EnKF. Introducing time dependence in the forecast error covariance is very intricate in standard 4D-Var. This has been one of the main motivations driving the flourishing of hybrid ensemble variational methods [23, 24]. In particular, when the background error covariance in the cost function is supplied by an ensemble, the formulation 4D-Var (En4D-Var) is obtained [25]. The En4D-Var retains flow-dependent covariances but optimizes the entire trajectory within each window rather than issuing instantaneous updates. The trade-off is computational cost as well as its high sensitivity to the algorithm initialization, in particular to the initial ensemble.
EnKF and nudging have already been applied to the Sabra shell model [26], yet prior studies never systematically varied the observational sparseness or the set of measured shells, nor did they analyze the implications of data assimilation for the energy cascade.
@@
In this work, we address these gaps by measuring different sets of shells at various observation sparseness, ranging from the slowest to the fastest time scales within the inertial range. This corresponds to passing from discrete to continuous sampling of inertial range dynamics. This setup allows us to assess how the temporal resolution, relative to the local dynamical time scale, influences the assimilation process (Fig. 1(b)). We show that measuring at rates faster than the turnover time of the observed scales, and covering at least two adjacent shells, enables full synchronization of the larger scales. Under these conditions, the EnKF delivers superior performance, provided that appropriate scale-aware inflation is applied to mitigate sampling errors. We also discuss the limitations of Nudging in propagating corrections across scales and highlight the computational trade-offs associated with the use of EnKF and En4D-Var.
@@
The remainder of the paper is organized as follows: In Section II, we introduce the shell model. Section III describes the EnKF algorithm, with special attention to the structure of the covariance. Section IV details the experimental setup, while Section V presents a comprehensive analysis of the EnKF results. Comparisons with Nudging and En4D-Var are discussed in Section VI and Section VII, respectively. Conclusions and future perspectives are summarized in Section VIII, and details on the numerical integration of the model are given in Appendix A while technical details on the inflation technique are discussed in Appendix B
@@
Figure 1: (a) Instantaneous estimate of characteristic turnover times. During one full oscillation of the slow shell $n=4$, the fast shell $n=15$—representative of the dissipative range—undergoes several hundred oscillations. (b) Quantitative visualization of the natural frequencies involved. The observation intervals examined are highlighted: $\Delta t_{\text{obs}}=\tau_{15}$ yields quasi-continuous inertial-range measurements, $\Delta t_{\text{obs}}=\tau_{9}$ represents an intermediate regime, and $\Delta t_{\text{obs}}=\tau_{4}$ corresponds to discrete sampling. (c) Normalized probability-density functions of shell amplitudes. Slow shells display nearly Gaussian statistics, whereas fast shells exhibit heavy-tailed, intermittent behavior.
@@ II Shell Models for Turbulence
-- $ck_{n-1}u_{n-1}u_{n-2})\,.$
The system of coupled differential equations is closed by imposing the boundary conditions $u_{-2}=u_{-1}=u_{N}=u_{N+1}=0$.
We have chosen to work with this model because the two-point correlation is diagonal $\langle u_{n}u_{m}^{*}\rangle=0$ for $n\neq m$, and the only vanishing three-point correlations are the consecutive ones $\langle u_{n-1}u_{n}^{*}u_{n+1}\rangle$. These features imply substantial analytical and numerical simplifications, enforce the locality of energy transfer, and arise from an inherent phase invariance under the transformation $u_{n}\rightarrow u_{n}e^{i\theta}$, provided that
@@
$S_{p}(k_{n})=\left\langle|u_{n}|^{p}\right\rangle_{T}\ ,$ (4)
where $\langle\cdot\rangle_{T}$ denotes a time average taken in the statistically stationary regime. In the inertial range they exhibit behavior $S_{p}(k_{n})\propto k_{n}^{-\zeta(p)}$, with $\zeta(p)$ anomalous scaling exponents that are remarkably close to those measured in Navier-Stokes turbulence [15, 12]. The power law dependence highlights the multiscale character of the energy cascade and the presence of multiple characteristic timescales (Fig. 1(b)), which can be estimated as
@@ III Ensemble Kalman Filter for systems without circular symmetry
In addition to being inherently non-linear, shell models exhibit scale-dependent statistics, being approximately Gaussian at large scales but increasingly non-Gaussian with heavy tails at smaller ones. This variability requires a filtering method that is robust to deviations from Gaussianity. Among the available quasi non-linear ensemble Kalman filtering approaches, we adopt the stochastic variant [28], which by introducing Gaussian perturbations during the correction step, mitigates the impact of outliers [29] and results particularly well-suited to the multi-scale statistical structure of shell models.
@@
$\hat{\bm{\Sigma}}=\begin{bmatrix}\Re(\hat{\bm{\Sigma}}_{\mathbb{C}}+\hat{\bm{\Sigma}}_{p})&-\Im(\hat{\bm{\Sigma}}_{\mathbb{C}}-\hat{\bm{\Sigma}}_{p})\\ \Im(\hat{\bm{\Sigma}}_{\mathbb{C}}+\hat{\bm{\Sigma}}_{p})&\ \ \Re(\hat{\bm{\Sigma}}_{\mathbb{C}}-\hat{\bm{\Sigma}}_{p})\end{bmatrix}\ ,$ (7)
@@
$+\!\bm{K}(t_{\text{obs}})\!\left[\bm{Z}(t_{\text{obs}})\!-\!\bm{H}\hat{\bm{U}}^{(j)}(t_{\text{obs}})\!-\!\bm{V}^{(j)}(t_{\text{obs}})\right]\,,$
@@
The posterior ensemble moves closer to the observations when $\bm{R}$ is small and remains near the prior when $\hat{\bm{\Sigma}}(t_{\text{obs}})$ is small. For Gaussian errors and linear observations and the forecast model, the Kalman gain optimally balances these contributions by minimizing the trace of posterior covariance $\tilde{\bm{\Sigma}}$, satisfying $\partial\,\mathrm{tr}(\tilde{\bm{\Sigma}})/\partial\bm{K}=0$ [16]. Stochastic perturbations $\bm{V}^{(j)}$ ensure that the updated ensemble remains a consistent sample of the filtering distribution: specifically, they avoid underestimating the “theoretical” posterior covariance [28].
@@
The posterior covariance is then obtained recursively as in the Kalman filter update according to $\tilde{\bm{\Sigma}}(t_{\text{obs}})=(\bm{I}-\bm{K}\bm{H})\hat{\bm{\Sigma}}(t_{\text{obs}})$, from where we see that its determinant is bound to be smaller or equal to one, reflecting a reduction in uncertainty. The posterior ensemble is obtained directly, member-by-member, by executing Eq. (9) (green circles in Fig. 2). The EnKF thus solves sequentially, likewise the Kalman filter [33], the state-estimation problem, by alternating the forecast (prior) and the analysis (posterior) steps, until the onset of the saturation regime. In this regime, the ensemble mean (under the assumption of Gaussian errors) has already converged near to the true state, and the ensemble spread stabilizes to a nearly constant value; at this point, the EnKF is no longer able to reduce uncertainty further, despite continuing to alternate between forecast and analysis steps.
Furthermore, when the dynamical model is free from errors (as in our setting) and deterministic, and under the aforementioned Gaussian case, this convergence guarantees complete synchronization with the ground-truth dynamics, as evidenced by the vanishing of the time- and ensemble-averaged error: $\langle(\bm{U}_{n}-\tilde{\bm{U}}_{n})^{2}\rangle_{T,L}=0$. A schematic illustration of the functioning of the EnKF is provided in Fig. 2.
As anticipated in the introduction, the limited size of the ensemble can lead to instabilities in the filter. These issues must be addressed using specific techniques, often referred to as “recipes”, such as localization and inflation [7, 8, 16]. The approach we adopted to mitigate these pathologies is described in detail in Appendix B.

### v2

@@ Abstract
We explore the potential of Data-Assimilation (DA) within the multi-scale framework of a shell model of turbulence, with a focus on the Ensemble Kalman Filter (EnKF). The central objective is to understand how measuring mesoscales (i.e., inertial-range scales) enhances the prediction of both large-scale and small-scale intermittent variables, by systematically varying observation frequency and the set of measured scales. We demonstrate that measurements conducted at frequencies that exceed those of the observed scales enable full synchronization of larger scales, provided that at least two adjacent mesoscale are measured. In addition, we benchmark the EnKF against two other DA methods, namely Nudging and Ensemble 4D-Var. EnKF is clearly superior to the former, and comparable with the latter but achieving the result with a lower computational complexity. Moreover, our results underscore the need for a tailored, scale-aware inflation technique to stabilize the assimilation process, preventing filter divergence and ensuring robust convergence.
@@ I Introduction
In order to overcome these limitations, the Data Assimilation (DA) framework was developed. Originally motivated by meteorological forecasting, DA provides systematic methods to combine model forecasts with sparse and noisy observations, improving the estimation of system states [6, 7, 8]. In chaotic systems, DA offers a practical strategy to counteract the exponential divergence of trajectories by acting as an external forcing—analogous to a drag—that competes with internal instabilities [9]. This competition naturally defines a synchronization threshold: When observational corrections outweigh the intrinsic divergence rate, synchronization can be achieved, and thus effective predictability.
@@
This work mainly focuses on investigating the performance of a cornerstone DA technique, the Ensemble Kalman Filter (EnKF) [10, 11], when applied to a shell model for turbulence [12, 13, 14, 15]. Our objective is to reconstruct the unobserved scales from sparse and noisy observations of the observed ones. Shell models emulate the turbulent energy cascade in a discretized Fourier space with logarithmically spaced wavenumbers, yielding several decades of scales and an equally wide spectrum of characteristic turnover times. Assimilating information measured at one scale to correct predictions at another is therefore an intrinsically multiscale challenge: updates must bridge slow, energy-containing modes and fast, highly intermittent ones.
@@
The EnKF advances an ensemble of model realizations and, at every assimilation instant, applies linear updates, provided the forecast and observation errors are Gaussian. Because the ensemble itself furnishes flow-dependent (prior) error covariances, the EnKF does not need to employ the tangent-linear and adjoint operators demanded by adjoint variational methods. However, in high-dimensional settings, computational constraints limit the ensemble size, introducing sampling noise that can erode covariance rank and trigger filter divergence [7, 16]. Practitioners therefore resort to inflation and covariance localization to maintain the need of ensemble spread and numerical stability [17, 18, 19, 20]. Despite these practical remedies, the EnKF and its extensions remain among the most widely used and readily implementable data assimilation methods for large-scale geophysical problems [8, 16]. These limitations become particularly important in the multiscale, strongly intermittent dynamics of turbulence. The shell model considered here spans a few decades of scales and turnover times, and it faithfully reproduces the heavy-tailed, non-Gaussian statistics characteristic of small-scale turbulent fluctuations (see Fig. 1(c)). We adopt the stochastic EnKF [11] , which injects controlled Gaussian perturbations into the observations at every analysis step. This added noise mitigates the underestimation of the analysis uncertainty due to sampling error, enhances ensemble dispersion, and partly compensates for the non-Gaussian heavy tails generated by the fastest, most intermittent scales, thereby improving filter robustness in the highly nonlinear, multiscale regime addressed here. The present contribution therefore constitutes a first step toward systematically tailoring ensemble-based DA techniques to highly turbulent flows.
@@
We shall benchmark the EnKF against two Data Assimilation (DA) methods at opposite ends of the DA complexity hierarchy. First, Nudging [21], which can be viewed as a stripped-down single-member analogue of the EnKF. By adding continuous relaxation terms toward observations, it requires almost no statistical assumptions and is computationally inexpensive, but its strictly local corrections struggle to propagate information to unmeasured scales. The propagation of information from observations throughout the full domain is done solely via the model dynamics, and, as opposed to the EnKF it does not provide a quantification of the uncertainty in the state estimate.
@@
Second, we consider the so-called four‐dimensional variational method (4D-Var) [22]. 4D-Var seeks the model trajectory that minimizes a cost function over a finite assimilation window, thereby distributing corrections smoothly in time. The 4D-Var is formulated as a nonlinear optimization problem, as opposed to the linear-like update in the EnKF. Introducing time dependence in the forecast error covariance is very intricate in standard 4D-Var. This has been one of the main motivations driving the flourishing of hybrid ensemble variational methods [23, 24]. In particular, when the background error covariance in the cost function is supplied fully or partly by an ensemble, the formulation En4D-Var is obtained [25]. The En4D-Var retains flow-dependent covariances but optimizes the entire trajectory within each window rather than issuing instantaneous updates. It is worth recalling that, in the linear case, 4D-Var and the Kalman filter are exactly equivalent and yield the same analysis at the end of the assimilation window, which makes their comparison particularly compelling in our strongly nonlinear model, where this equivalence no longer holds.
Data assimilation–based reconstruction from spectrally restricted observations has already been explored, by applying 4D-Var to a barotropic $\beta$-plane model to recover small-scale features [26], and by applying EnKF and nudging to the Sabra shell model [27]. Futhermore, in the context of coupled DA, recent works have focused on adapting the EnKF to perform coupled atmospheric-ocean DA under various spatial and temporal scale separations and observational strength [28, 29]. However, none of these studies have systematically varied the observational sparsity or the subset of observed scales, analyzing in detail the implications of DA on the energy cascade process.
@@
In this work, we address these gaps by measuring different sets of shells at various observation sparseness, ranging from the slowest to the fastest time scales within the inertial range. This corresponds to passing from discrete to continuous sampling of inertial range dynamics. This setup allows us to assess how the temporal resolution, relative to the local dynamical time scale, influences the assimilation process. We show that measuring at rates faster than the turnover time of the observed scales, and covering at least two adjacent shells, enables full synchronization of the larger scales. Under these conditions, the EnKF delivers superior performance, provided that appropriate scale-aware inflation is applied to mitigate sampling errors. We also discuss the limitations of Nudging in propagating corrections across scales and highlight the computational trade-offs associated with the use of EnKF and En4D-Var. We focus on a data assimilation setting in which both large and small scales are inferred from relatively sparse measurements at intermediate scales; although this is not the only possible configuration (for instance, one could directly observe the largest scales), it reflects the typical situation encountered in turbulent flows, where many degrees of freedom cooperate to sustain strongly out-of-equilibrium dynamics and no clear scale separation is available, in contrast to multiscale extensions of the Lorenz–96 model [30] where a slow–fast splitting is imposed a priori.
@@
The remainder of the paper is organized as follows: In Section II, we introduce the shell model. Section III describes the EnKF algorithm, with special attention to the structure of the covariance and to the proposed inflation technique. Section IV details the experimental setup, while Section V presents a comprehensive analysis of the EnKF results. Comparisons with Nudging and En4D-Var are discussed in Section VI and Section VII, respectively. Conclusions and future perspectives are summarized in Section VIII, and details on the numerical integration of the model are given in Appendix A.
@@
Figure 1: Multi-scale character of the Sabra Model (a) Instantaneous estimate of characteristic turnover times. During one full oscillation of the slow shell $n=4$, the fast shell n=15n=15—representative of the dissipative range—undergoes several hundred oscillations. (b) Quantitative visualization of the natural frequencies involved. The observation intervals examined are highlighted: $\Delta t_{\text{obs}}=\tau_{15}$ yields quasi-continuous inertial-range measurements, $\Delta t_{\text{obs}}=\tau_{9}$ represents an intermediate regime, and $\Delta t_{\text{obs}}=\tau_{4}$ corresponds to discrete sampling. (c) Normalized probability-density functions of shell amplitudes. Slow shells display nearly Gaussian statistics, whereas fast shells exhibit heavy-tailed, intermittent behavior.
@@ II Shell Models for Turbulence
- $ck_{n-1}u_{n-1}u_{n-2})\,$,
where * denotes the complex conjugation. The system of coupled differential equations is closed by imposing the boundary conditions $u_{-2}=u_{-1}=u_{N}=u_{N+1}=0$.
We have chosen to work with this model because the two-point correlation is diagonal $\langle u_{n}u_{m}^{*}\rangle_{T}=0$ for $n\neq m$, and the only vanishing three-point correlations are the consecutive ones $\langle u_{n-1}u_{n}^{*}u_{n+1}\rangle_{T}$ (where $\langle\cdot\rangle_{T}$ denotes a time average taken in the statistically stationary regime). These features imply substantial analytical and numerical simplifications, enforce the locality of energy transfer, and arise from an inherent phase invariance under the transformation $u_{n}\rightarrow u_{n}e^{i\theta}$, provided that
@@
$S_{p}(k_{n})=\left\langle|u_{n}|^{p}\right\rangle_{T}\ .$ (4)
In the inertial range they exhibit behavior $S_{p}(k_{n})\propto k_{n}^{-\zeta(p)}$, with the exponents $\zeta(p)$ remarkably close to those measured in Navier-Stokes turbulence [15, 32], but “anomalous” in the sense they deviate from the Kolmogorov prediction (1941) [12] $\zeta(p)=p/3$. The power law dependence highlights the multiscale character of the energy cascade and the presence of multiple characteristic timescales (Fig. 1(a) and (b)), which can be estimated as
@@
We close this section mentioning that even though shell models may seem a simplification of 3D turbulent flows, in particular for the absence of a spatial structures, they preserve the multiscale character of turbulence with the presence of many characteristic times, which is one of the main difficulty when facing with predictability problems in turbulent flows. They represent a unique test bed where DA tools can be checked on fluid-like systems spanning many decades in frequencies and with strong non-Gaussians fluctuations (see Fig. 1(c)).
@@ III Ensemble Kalman Filter for systems without circular symmetry
In addition to being inherently non-linear, shell models exhibit scale-dependent statistics, being approximately Gaussian at large scales but increasingly non-Gaussian with heavy tails at smaller ones. This variability requires a filtering method that is robust to deviations from Gaussianity. Among the available weakly non-linear ensemble Kalman filtering approaches, we adopt the stochastic variant in which each ensemble member assimilates a perturbed version of the observations [33]. In the stochastic EnKF, perturbations for the observations are drawn from the same Gaussian distribution assumed to describe the observational error. This choice mitigates the collapse of the analysis members onto the observations with consequent reduction of spread, the underestimation of the posterior covariance [33, 8], and the impact of outliers [34]. We thus consider the stochastic EnKF to be particularly well-suited to the multi-scale statistical structure of shell models.
@@
[…]
@@
$\hskip-50.00008pt+\!\bm{K}(t_{\text{obs}})\!\left[\bm{Z}(t_{\text{obs}})\!-\!\bm{H}\hat{\bm{U}}^{(j)}(t_{\text{obs}})\!-\!\bm{V}^{(j)}(t_{\text{obs}})\right]\,$,
@@
The posterior ensemble moves closer to the observations when $\bm{R}$ is small and remains near the prior when $\hat{\bm{\Sigma}}(t_{\text{obs}})$ is small. For Gaussian errors and linear observations of the forecast model, the Kalman gain optimally balances these contributions by minimizing the trace of posterior covariance $\tilde{\bm{\Sigma}}$, satisfying $\partial\,\mathrm{tr}(\tilde{\bm{\Sigma}})/\partial\bm{K}=0$ [16]. Stochastic perturbations $\bm{V}^{(j)}$ ensure that the updated ensemble remains a consistent sample of the filtering distribution: specifically, they avoid underestimating the “theoretical” posterior covariance [33].
@@
The posterior covariance is then obtained recursively as in the Kalman filter update according to $\tilde{\bm{\Sigma}}(t_{\text{obs}})=(\bm{I}-\bm{K}\bm{H})\hat{\bm{\Sigma}}(t_{\text{obs}})$, from where we see that its determinant is bound to be smaller or equal to one, reflecting a reduction of uncertainty. The posterior ensemble is obtained directly, member-by-member, by executing Eq. (9) (green circles in Fig. 2). The EnKF thus solves sequentially, likewise the Kalman filter [38], the state-estimation problem, by alternating the forecast (prior) and the analysis (posterior) steps, until the onset of the saturation regime, where the ensemble mean has likely converged near to the true state. A schematic illustration of the functioning of the EnKF is provided in Fig. 2.
III.1 Scale Aware Inflation Strategy
As a Monte Carlo-based method, EnKF is inherently prone to instability issues due to sampling errors. These arise from the inevitable limited number of ensemble members in comparison to what is needed to properly describe the system’s effective degrees of freedom [7]. For Gaussian and linear systems, the skill of stochastic EnKF monotonically converges to the ideal KF in the limit of an infinitely large ensemble. The minimum number of members required for satisfactory performance and computational affordability depends on the model-and- observational scenarios, as well as on the type of EnKF. For chaotic dynamics, in principle, there is a relation with the number of unstable modes of dynamics [9]. When the ensemble size is small, the estimated covariance matrix tends to contain spurious correlations, typically manifesting as underestimated variances and overestimated cross-correlations at large distance. While the latter are usually understood in physical space, the same is often the case also in spectral domain, whereby the actual correlation between scales/frequencies decreases with the frequency gap, while the ensemble-based correlations do not decrease similarly fast or not at all. These artificial, non-physical correlations pose a serious problem: when a new observation is assimilated, the update step may inadvertently modify variables that should remain unaffected, excessively reducing their associated error variance. In the next assimilation cycle, the ensemble spread becomes so narrow that even small, incorrect adjustments can propagate across unrelated variables. The EnKF assigns insufficient weight to incoming observations, leading to filter divergence, which is particularly problematic in chaotic systems [39]. The resulting error perturbations increase the stiffness of the system and can ultimately cause catastrophic divergence, a finite-time numerical blow-up of the forecast [17]. The indisputable success of EnKF methods in high-dimension chaotic systems is also related to the development of a plethora of approaches, and ad-hoc fixes, to counteract the effect of the sampling error. These efforts can all be framed under two algorithmic solutions: localization and inflation [7, 8, 16]. Localization consists in restricting the effect of the analysis update to the vicinity (or local domain) of each observation and is a widely used and often indispensable technique in medium- to large-scale systems. The size of the localization domain is usually chosen based on practical arguments and may require numerical tuning, although recent studies have introduced scale-dependent localization approaches [40, 41]. In the shell model, the energy cascade is inherently local, with interactions occurring primarily between neighboring scales. This makes it natural to impose a tridiagonal structure on the covariance matrix as a form of spectral localization. However, our experiments reveal that such localization—either in the form of a tridiagonal approximation or of a stricter spectral correlation cutoff—while successful in eliminating spurious correlations due to sampling error, leads to physically imbalanced analyses causing larger forecast errors. We therefore omit using localization and instead shall demonstrate that an appropriate inflation strategy will suffice. We introduce a novel scale-aware inflation approach that mitigates the effects of undersampling without sacrificing physical consistency. Inflation consists of artificially increasing the ensemble-based error covariance [19], and is usually performed in two alternative ways, additive or multiplicative [7]. It can be applied to either the forecast (prior) or the analysis (posterior) error covariance, with different effects. In additive inflation, a random term is added to each ensemble member to maintain the diversity and the spread of the ensemble. In multiplicative inflation, the ensemble-based error covariance, either prior or posterior, is inflated by a scalar factor. This inflates the ensemble variance by rescaling each posterior ensemble member around the mean as:
$g_{n}(t)\tilde{U}_{n}+(1-g_{n}(t))\tilde{\mu}_{n}\to\tilde{U}_{n}\ ,$ (12)
where $\tilde{\mu}_{n}$ is the posterior mean and $g_{n}(t)\geq 1$ is the inflation coefficient. This transformation increases the dispersion of the ensemble around the mean and rescales the covariance element $\tilde{\Sigma}_{nm}$ to $g_{n}g_{m}\tilde{\Sigma}_{nm}$. Sacher and Bartello [18] showed that the sampling error in the Kalman gain estimate tends to scale directly proportional to the magnitude of the gain itself. This implies that larger corrections from observations require stronger inflation to maintain stability. It is therefore desirable to inflate the ensemble more in regions where observations exert a stronger influence. Motivated by this, we propose a scale-aware inflation scheme, where the spread reduction is compensated by:
with $\hat{\Sigma}_{nn}$ and $\tilde{\Sigma}_{nn}$ denoting the prior and posterior variances, respectively, and $\lambda\in{\mathbb{R}}^{+}$ the inflation strength, which is a hyperparameter to be tuned. This formulation is inspired by the original work of Zhang et al. [42], which used a similar expression but normalized by the posterior variance $\tilde{\Sigma}_{nn}$ instead of $\hat{\Sigma}_{nn}$}. Since, statistically, $\langle\hat{\Sigma}_{nn}\rangle_{T}\geq\langle\tilde{\Sigma}_{nn}\rangle_{T}$, our version yields a smaller inflation coefficient for the same spread variation and $\lambda$, resulting in a less aggressive inflation. The choice of the parameter $\lambda$ is dependent on the system and must be carefully tuned to ensure effective assimilation. Although our inflation strategy is scale-sensitive, tuning $\lambda$ adds computational cost. A detailed analysis of its effect on analysis statistics, filter stability, and comparison with other techniques lies beyond the scope of this work. However, even with extensive adjustment of $\lambda$, the proposed inflation strategy turns out to be very successful. It has been effective in stabilizing the assimilation process and in substantially reducing the minimum ensemble size needed to achieve satisfactorily performance and avoid filter divergence. Further and dedicated developments of the proposed scale-aware inflation method will be the content of follow-up works.
Numerical details on inflation prefactor $\lambda$ used are provided in Appendix B.